a pencil and a rubber cost #54 together.Four pencils cost the same as five rubbers. Find the cost of each

Answer by mananth(16946)

a pencil and a rubber cost Rs 54 together.Four pencils cost the same as five rubbers. Find the cost of each
let cost of pencil be x
and rubber be y
x+y=54
Four pencils cost the same as five rubbers.
4x=5y
x=5y/4
substitute x in First equation
x+y=54
5y/4 +y=54
multiply all the terms by 4
5y+4y=216
9y=216
y=24
You find out x
from equation x+y=54